OpenStudy (anonymous):

Translate the sentence into an equation. 1. Nine less than the product of three and the number x is equal to one-half the sum of x and 12. 2. Three times the sum of a and b is equal to five times c. 3. The number x is divided by the number y is the same as six less than three times the difference of x and y.

OpenStudy (anonymous):

1. 9 - (3x) = (x+12)/2 2. 3(a+b) = 5c 3. x/y = 6-(3(x-y))